If refinishing your basement at a cost of $45,729 would add $21,688 to the resale value of your home, what percent of your cost is recouped?
Answer by Mathtut(3670)

cost is 45729
:
21688 is what percent of 45729
:
21688=(x)(45729)
:
x=21688/45729=.4743(100)=47.43%
:
so 47.43 % of the cost was recouped with the value increase